# On the PM2.5 – Mortality Relationship: A Bayesian Model for Spatio-Temporal Confounding

### v2

@@ Appendix D Details on Competing Models
    $D_{jt}=\begin{cases}1&\mbox{if the day $t$ is in the $j$-th season}\\ 0&\mbox{otherwise}\end{cases}$
@@
  • JDZ: the model proposed by Janes et al. (2007), where the local exposure effect is of interest, hence $\mathbf{A}_{it}=(X_{it}-\widehat{X}_{it})^{\prime}$ and $\mathbf{C}_{it}=(1,\widehat{X}_{it},{{\bf M}}_{it}^{\prime},\mathbf{B}_{t}^{*\prime})^{\prime}$. Here, $\widehat{X}_{it}$ denotes the overall temporal trend (see also Section 3), and $\mathbf{B}_{t}^{*}$ denotes the vector of basis functions from a natural cubic spline with $df-1$ degrees of freedom, evaluated at time $t$.
@@ E.1 Data Generation Mechanism
To simulate the variables, we use the following data generation mechanism. The spatial domain comprises ${N}=117$ contiguous HDs, while the temporal dimension is set to ${T}=600$. Let ${\bf X}=({\bf X}_{1},\dots,{\bf X}_{{T}})^{\prime}$, with ${\bf X}_{t}=(X_{1t},\dots,X_{{N}t})^{\prime}$, $t=1,\dots,{T}$, be the exposure, and let ${\bf Z}=({\bf Z}_{1},\dots,{\bf Z}_{{T}})^{\prime}$, with ${\bf Z}_{t}=(Z_{1t},\dots,Z_{{N}t})^{\prime}$, $t=1,\dots,{T}$, be the unmeasured confounder. It is thus assumed that
$\begin{bmatrix}\mathbf{X}\\ \mathbf{Z}\end{bmatrix}\sim\mathcal{N}\left(\begin{bmatrix}\bm{\mu}_{x}\\ \bm{\mu}_{z}\end{bmatrix},\begin{bmatrix}\bm{\varSigma}_{x}&\bm{\varSigma}_{xz}\\ \bm{\varSigma}_{zx}&\bm{\varSigma}_{z}\end{bmatrix}\right)\,,$ (11)
@@
$\vartheta_{it}$ $={\color[rgb]{0,0,0}o_{it}+\beta_{0}+\beta_{1it}X_{it}+Z_{it}+u_{it},\quad u_{it}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\tau_{u}^{2})}\,,$ (14)
@@ E.2 Performance Assessment
Tables E.4E.6 indicate that, SDGLMC is the best model among those compared in terms of $MSE_{avg}$, $MAB$, and $MAE_{avg}$ for all scenarios. Recalling the decomposition of $\beta_{1it}$ in Section 3.1 of the main manuscript, we have also computed the average mean squared error for each of the components of the exposure effect – global, temporal and spatial. SDGLMC provides the lowest values in most of the cases, and the improvements over the other approaches are more pronounced for the temporal component.
@@
Global $100(\exp\{10\overline{\delta}_{1}\}-1)$ 0.115 34.568 0.193 0.142 0.225 0.275
@@
Global $100(\exp\{10\overline{\delta}_{1}\}-1)$ 5.246 13.331 4.807 4.897 6.568 4.893
@@
Global $100(\exp\{10\overline{\delta}_{1}\}-1)$ 0.509 29.444 0.370 0.310 0.423 0.417
@@
Spatial $100(\exp\{10\delta_{1i}\}-1)$ 0.216 0.385 0.498 0.500 0.506 0.502
Table E.5: Simulation Study 1: $MAB$ calculated for each model considered under each Scenario. The lowest values are indicated in italic.
@@
Global $100(\exp\{10\overline{\delta}_{1}\}-1)$ 0.295 5.856 0.189 0.022 0.340 0.316
@@
Global $100(\exp\{10\overline{\delta}_{1}\}-1)$ 2.280 3.647 2.183 2.204 2.555 2.203
@@
Global $100(\exp\{10\overline{\delta}_{1}\}-1)$ 0.599 5.413 0.479 0.403 0.529 0.523
@@
Global $100(\exp\{10\overline{\delta}_{1}\}-1)$ 0.297 5.856 0.340 0.286 0.400 0.422
@@
Global $100(\exp\{10\overline{\delta}_{1}\}-1)$ 2.280 3.647 2.183 2.204 2.555 2.203
@@
Global $100(\exp\{10\overline{\delta}_{1}\}-1)$ 0.599 5.413 0.479 0.404 0.529 0.523
@@ Appendix F Simulation Study 2
F.1 On The Removal of Large-Scale Temporal Trend.
We argue that the term $X_{it}-\widehat{X}_{it}$, for $i=1,\dots,{N}$ and $t=1,\dots,{T}$, should be used in place of simply $X_{it}$ in the DGLM. Removing the overall temporal trend from the exposure not only follows from theoretical arguments (see Equation (3) in the main paper), but also it is crucial to control for confounding bias. The local variations in the exposure are in fact less likely to be confounded by the unmeasured confounder, which is assumed to be spatially and temporally smooth (Janes et al., 2007, Greven et al., 2011) – see also Section 3 of the manuscript. To illustrate this with simulated data, we compare the proposed SDGLMC with a reparameterized model that does not remove the overall trend from the exposure. Essentially, if the first-order Taylor’s series is evaluated at 00, Equation (6) in the paper can be rewritten as:
${\color[rgb]{0,0,0}\vartheta_{it}=o_{it}+\beta_{0it}^{*}+\beta_{1t}^{*}X_{it}+{\mathbf{M}}_{it}^{\prime}\bm{\alpha}+\tilde{u}_{it}\,,}$
@@
We refer to this model as SDGLMC_single, since it is a single-step approach that does not require to estimate the overall spatio-temporal trend.
@@ G.1 Comparison with Restricted Model
${\color[rgb]{0,0,0}\bm{\vartheta}_{t}=\mathbf{o}_{t}+{\bf 1}_{N}(\overline{\delta}_{0}+\delta_{0t})+\tilde{\bm{\delta}}_{0}+\bm{\delta}_{0t}^{*}+\mathbf{\tilde{X}}_{t}(\overline{\delta}_{1}+\delta_{1t})+\mathbf{\tilde{X}}_{t}\odot\tilde{\bm{\delta}}_{1}+\tilde{{\bf u}}_{t}\,.}$
@@
However, it is worth noting that SDGLMCX provides greater estimates than SDGLMC during the summer, while during the winter the estimates are quite similar. This suggests that the confounding effect may vary across time. We hypothesize that this is due to air temperature exceeding a certain threshold; in other words, confounding may occur due to heat waves. In fact, heat waves are demonstrated to have an increased effect on mortality during summer days with relatively high levels of PM2.5 concentrations, and this positive correlation may be the source of the overestimation (Analitis et al., 2008, 2014). Another study observed a similar result for the population aged 75 years or more (Analitis et al., 2018). Another possible explanation for the overestimation could be the modification of the PM2.5 chemical composition, with high levels of the most toxic compounds developing in summer (Vinson et al., 2020).

### v3

@@ Appendix D Details on Competing Models
[…]
@@
  • JDZ: the model proposed by Janes et al. (2007), where the local exposure effect is of interest, hence $\mathbf{A}_{it}=(X_{it}-\widehat{X}_{it})^{\prime}$ and $\mathbf{C}_{it}=(1,\widehat{X}_{it},{{\bf M}}_{it}^{\prime},\mathbf{B}_{t}^{*\prime})^{\prime}$. Here, $\widehat{X}_{it}$ denotes the overall temporal trend (see also Section 3), and $\mathbf{B}_{t}^{*}$ denotes the vector of basis functions from a natural cubic spline with df1df-1 degrees of freedom, evaluated at time $t$.
@@ E.1 Data Generation Mechanism
To simulate the variables, we use the following data generation mechanism. The spatial domain comprises N=117{N}=117 contiguous HDs, while the temporal dimension is set to T=600{T}=600. Let ${\bf X}=({\bf X}_{1},\dots,{\bf X}_{{T}})^{\prime}$, with ${\bf X}_{t}=(X_{1t},\dots,X_{{N}t})^{\prime}$, $t=1,\dots,{T}$, be the exposure, and let ${\bf Z}=({\bf Z}_{1},\dots,{\bf Z}_{{T}})^{\prime}$, with ${\bf Z}_{t}=(Z_{1t},\dots,Z_{{N}t})^{\prime}$, $t=1,\dots,{T}$, be the unmeasured confounder. It is thus assumed that
[…]
@@
$\vartheta_{it}$ $=o_{it}+\beta_{0}+\beta_{1it}X_{it}+Z_{it}+u_{it},\quad u_{it}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\tau_{u}^{2})\,$, (13)
@@ E.2 Performance Assessment
Tables E.4E.6 indicate that, SDGLMC is the best model among those compared in terms of $MSE_{avg}$, $MAB$, and $MAE_{avg}$ for all scenarios. Recalling the decomposition of $\beta_{1it}$ in Section 3.1 of the main manuscript, we have also computed the average mean squared error for each of the components of the exposure effect – baseline, temporal and spatial. SDGLMC provides the lowest values in most of the cases, and the improvements over the other approaches are more pronounced for the temporal component.
@@
Baseline $100(\exp\{10\overline{\delta}_{1}\}-1)$ 0.115 34.568 0.193 0.142 0.225 0.275
@@
Baseline $100(\exp\{10\overline{\delta}_{1}\}-1)$ 5.246 13.331 4.807 4.897 6.568 4.893
@@
Baseline $100(\exp\{10\overline{\delta}_{1}\}-1)$ 0.509 29.444 0.370 0.310 0.423 0.417
@@
Spatial $100(\exp\{10\tilde{\delta}_{1i}\}-1)$ 0.216 0.385 0.498 0.500 0.506 0.502
Table E.5: Simulation Study 1: MABMAB calculated for each model considered under each Scenario. The lowest values are indicated in italic.
@@
Baseline $100(\exp\{10\overline{\delta}_{1}\}-1)$ 0.295 5.856 0.189 0.022 0.340 0.316
@@
Baseline $100(\exp\{10\overline{\delta}_{1}\}-1)$ 2.280 3.647 2.183 2.204 2.555 2.203
@@
Baseline $100(\exp\{10\overline{\delta}_{1}\}-1)$ 0.599 5.413 0.479 0.403 0.529 0.523
@@
Baseline $100(\exp\{10\overline{\delta}_{1}\}-1)$ 0.297 5.856 0.340 0.286 0.400 0.422
@@
Baseline $100(\exp\{10\overline{\delta}_{1}\}-1)$ 2.280 3.647 2.183 2.204 2.555 2.203
@@
Baseline $100(\exp\{10\overline{\delta}_{1}\}-1)$ 0.599 5.413 0.479 0.404 0.529 0.523
@@ Appendix F Simulation Study 2
F.1 On The Removal of Large-Scale Spatio-Temporal Trend.
We argue that the term $X_{it}-\widehat{X}_{it}$, for $i=1,\dots,{N}$ and $t=1,\dots,{T}$, should be used in place of simply $X_{it}$ in the DGLM. Removing the large-scale spatio-temporal trend from the exposure not only follows from theoretical arguments (see Equation (3) in the main paper), but also it is crucial to control for confounding bias. The local variations in the exposure are in fact less likely to be confounded by the unmeasured confounder, which is assumed to be spatially and temporally smooth (Janes et al., 2007, Greven et al., 2011) – see also Section 3 of the manuscript. To illustrate this with simulated data, we compare the proposed SDGLMC with a reparameterized model that does not remove the large-scale trend from the exposure. Essentially, if the first-order Taylor’s series is evaluated at 0, Equation (5) in the paper can be rewritten as:
$\vartheta_{it}=o_{it}+\beta_{0it}^{*}+\beta_{1t}^{*}X_{it}+{\mathbf{M}}_{it}^{\prime}\bm{\alpha}+\tilde{u}_{it}\,$,
@@
We refer to this model as SDGLMC_single, since it is a single-step approach that does not require to estimate the large-scale spatio-temporal trend.
@@ G.1 Comparison with Restricted Model
$\bm{\vartheta}_{t}=\mathbf{o}_{t}+{\bf 1}_{N}(\overline{\delta}_{0}+\delta_{0t})+\tilde{\bm{\delta}}_{0}+\bm{\delta}_{0t}^{*}+\mathbf{\tilde{X}}_{t}(\overline{\delta}_{1}+\delta_{1t})+\mathbf{\tilde{X}}_{t}\odot\tilde{\bm{\delta}}_{1}+\tilde{{\bf u}}_{t}\,.$
@@
However, it is worth noting that SDGLMCX provides greater estimates than SDGLMC during the summer, while during the winter the estimates are quite similar. This suggests that the confounding effect may change over time. We hypothesize that this is due to air temperature exceeding a certain threshold; in other words, confounding may occur due to heat waves. In fact, heat waves are demonstrated to have an increased effect on mortality during summer days with relatively high levels of PM2.5 concentrations, and this positive correlation may be the source of the overestimation (Analitis et al., 2008, 2014). Another study observed a similar result for the population aged 75 years or more (Analitis et al., 2018). Another possible explanation for the overestimation could be the modification of the PM2.5 chemical composition, with high levels of the most toxic compounds developing in summer (Vinson et al., 2020).